Home‎ > ‎Crew 164‎ > ‎

0302-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 10
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title: Ready to Lose Sight of the Shore and Discover New Oceans!
Mission Status: The overall operations mission performance is nominal and routine operations are on-going.

Sol Activity Summary: EVA#10. Photo shooting session with flags. Food inventory. Maintenance and repair work on backpacks.

Look Ahead Plan: Continuation of routine mission operations and work on experiments: Two EVA's planned tomorrow, including a short one in the afternoon to redo the RMR Geology Experiment at correct coordinates.

Anomalies in work: Water leak observed close to the HAB entrance. ATV 4 gearbox issue. Observatory shutter closing issue (see Astronomy report).

Weather: Sunny and very windy.

Crew Physical Status: Some nose bleeding -due to dry atmosphere inside the HAB- has been reported. See HSO report for more details. 

EVA: EVA in the morning to explore canyons East of the HAB. The crew really enjoyed the view and scouted an interesting place to test the CRV.

Reports to be filed: 
- Commander report
- Engineering report
- Journalist report 
- EVA #10 report
- EVA #11A request
- EVA #11B request
- HSO report
- Astronomy report

Support Requested: /